### v1

# Self-Ensemble: Mitigating Confidence Distortion for Large Language Models

Abstract
Although Large Language Models (LLMs) perform well in general fields, they exhibit a confidence distortion problem on multi-choice question-answering (MCQA), particularly as the number of answer choices increases. Specifically, on MCQA with many choices, LLMs suffer from under-confidence in correct predictions and over-confidence in incorrect ones, leading to a substantially degraded performance. To solve this problem, we propose Self-Ensemble in this work. Our method splits the choices into several groups and ensembles LLM predictions across these groups to reach a final decision. The advantage of Self-Ensemble is its plug-and-play nature, where it can be integrated into existing LLM architecture based on a designed attention mask and positional encoding, without requiring labeled datasets for parameter tuning. Experimental results on three LLMs and datasets demonstrate that Self-Ensemble comprehensively addresses the confidence distortion problem of LLMs, outperforming standard inference as well as baseline methods.
Self-Ensemble: Mitigating Confidence Distortion for
Large Language Models
Zicheng Xu1,3thanks: Equal contribution, ordered by rolling dices., Guanchu Wang211footnotemark: 1, Guangyao Zheng3, Yu-Neng Chuang1, Alexander Szalay3, Xia Hu1, Vladimir Braverman3 1Rice University, 2University of North Carolina at Charlotte, 3Johns Hopkins University Correspondence: vova@cs.jhu.edu
1 Introduction
Large Language Models (LLMs) have exhibited remarkable performance in processing natural language information, such as the LLaMA Touvron et al. (2023), Mistral Jiang et al. (2024), and Deepseek Guo et al. (2025). Among various NLP tasks, multi-choice question-answering (MCQA) stands out as a standard and challenging benchmark to evaluate the reliability and reasoning ability of LLMs Wang et al. (2024c). It can significantly reduce the hallucinations by limiting the answer to a predefined set of choice aligned with human knowledge Anjum et al. (2024); Neeley et al. (2025).
While advanced LLMs perform well on standard MCQA benchmarks, such as MMLU Hendrycks et al. (2021), MathQA Amini et al. (2019), and GSM-8K Cobbe et al. (2021), they still face challenges when it comes to numerous choices that are closely related Wang et al. (2024a). In particular, we identify a critical problem: LLMs suffer from confidence distortion on MCQA tasks based on a comprehensive benchmark analysis. Specifically, LLM’s confidence in the correct choice tends to degrade, while its confidence in incorrect choices increases as the number of choices grows. This problem grows even more pronounced in the presence of additional noisy or partially relevant choices, thereby increasing the unreliable generations and erroneous predictions.
Figure 1: Self-Ensemble’s comprehensive performance on the QASC, TruthfulQA, and MMLU-Pro Biology datasets compared with baseline methods.
(a) LLaMA-3-8B (Correct)
(b) Mistral-7B-v0.1 (Correct)
(c) Qwen2-7B (Correct)
(d) LLaMA-3-8B (Incorrect)
(e) Mistral-7B-v0.1 (Incorrect)
(f) Qwen2-7B (Incorrect)
Figure 2: Proportion of model prediction probability exceeding a threshold on the QASC dataset, for each model under both correct- and incorrect-answer conditions.
To calibrate LLM predictions on MCQA, existing work focuses on post-processing LLM predictions to better align confidence with correctness. One common method is the vector scaling Guo et al. (2017). It adjusts the model outputs by applying a learned scaling vector and bias vector to the logits before applying the softmax function, which can smooth or sharpen the probabilities to better align with correctness. Another approach is Dirichlet Calibration Zong et al. (2024). It generalizes LLM’s prediction probability on each choice using a Dirichlet distribution to maximize the likelihood function on a validation set. However, a major limitation of these methods is their reliance on a labeled validation dataset to optimize calibration parameters. In practical scenarios, collecting high-quality validation data can be both labor-intensive and costly, which limits these techniques.
To overcome this limitation, we introduce Self-Ensemble for LLM calibration in this work. Unlike existing work, Self-Ensemble calibrates LLM predictions without relying on a labeled validation dataset for parameter tuning. Specifically, Self-Ensemble mitigates the confidence distortion problem through a divide-and-conquer process: it splits the choices into several groups of choice subsets; then ensembles LLM predictions on each group to achieve the final decision. To plug into existing LLM architecture, we design attention masking and positional re-encoding mechanisms that enable the divide-and-conquer process to be executed internally during inference. Figure 1 shows the comprehensive performance of Self-Ensemble on three datasets and LLMs, where our method shows significant improvement over standard inference and baseline methods by addressing the confidence distortion problem. To summarize, the contributions of this work are as follows:
  • Confidence distortion. We identify the confidence distortion problem of LLMs in MCQA task, where LLMs have under-confident problems on correct answers and over-confident problems on incorrect answers.
  • Self-Ensemble. We introduce Self-Ensemble to calibrate LLM confidence in MCQA task. We design attention masking and positional re-encoding to plug Self-Ensemble into existing LLM architectures.
  • Evaluation. Experimental results demonstrate that Self-Ensemble can comprehensively mitigate the under-confident problems on correct answers and over-confident problems on incorrect answers. It can effectively improve accuracy on benchmark datasets by 8% on average.
2 Preliminary
2.1 Notations
We consider an LLM $f$ on MCQA tasks in this work. We define an MCQA task with $K$ answer choices as $K$-choice QA, where the set of options is denoted as $\mathcal{C}\!=\!\{op_{1},op_{2},\cdots,op_{K}\}$. In this work, we refer to the setting with $K\!\leq\!4$ as few-choice QA, and with $K\!\geq\!8$ as many-choice QA. Our goal is to solve the degradation of LLM’s performance and confidence on the task of many-choice QA.
2.2 LLMs for MCQA
The two most effective ways for LLMs doing MCQA are verbal generation and token probability. In the verbal generation approach, the LLM generates the final answer as text outputs, corresponding to one of the choices. However, a key drawback is that it may fail to follow the expected output format, leading to answer extraction errors. In contrast, the token probability approach directly computes the probability of each choice token at the model’s final layer, selecting the answer with the highest probability. It offers a more reliable evaluation result and often reflects a model’s best achievable performance. In this work, we consider the token probability approach for MCQA.
Figure 3: LLMs ignore the correct choice and pick the incorrect one in the many-choice setting.
Model 2-choice 4-choice 6-choice 8-choice
LLaMA-3-8B 91.90 86.29 76.35 73.22
Mistral-7B-v0.1 86.83 75.81 65.98 57.78
Qwen2-7B 90.82 83.80 73.97 69.44
Average 89.85 81.97 72.10 66.81
Table 1: Accuracy of LLMs on the QASC dataset with different choice numbers.
Figure 4: Example of Self-Ensemble process on 4-choice QA.
3 Confidence Distortion and Self-Ensemble
We demonstrate the confidence distortion problem for LLMs and introduce our Self-Ensemble method to overcome this problem.
3.1 Confidence Distortion on Many-choice QA
We give Figure 2 to demonstrate the confidence distortion problems of LLMs on many-choice QA, where the experiment is on the QASC datasets Khot et al. (2020). Specifically, Figure 2 shows the proportion of $f([Q,\mathcal{C}])\!>\!\tau$ versus $\tau$ for $10\!<\!\tau\!<\!$ 1, where $f([Q,\mathcal{C}])\!>\!\tau$ represents LLM’s predicted probability on the correct or incorrect choice exceeds a threshold $\tau$. A high value of $f([Q,\mathcal{C}])\!>\!\tau$ proportion across $10\!\leq\!\tau\!\leq\!$ 1 indicates strong model confidence. To show the confidence distortion problem, we compare the $f([Q,\mathcal{C}])\!>\!\tau$ proportion across different choice settings with $K=2,4,6,8$ in Figure 2. On the correctly answered questions, the LLM achieves the highest $f([Q,\mathcal{C}])\!>\!\tau$ proportion when $K\!=\!2$, showing that the model becomes under-confident in many-choice scenarios. In contrast, for the incorrectly answered questions, the LLM shows the highest $f([Q,\mathcal{C}])\!>\!\tau$ proportion when $K\!=\!8$, indicating an over-confidence issue in many-choice scenarios.
The confidence distortion problem causes a degradation of LLM performance on many-choice QA. As shown in Figure 3, for the LLaMA-3-8B model, it can originally solve the 4-choice QA. However, when the same question is extended to 8 choices, while keeping the correct answer unchanged, the model selects an incorrect option. This failure is due to a decrease in confidence in the correct answer and an increase in confidence in incorrect options. We show a comprehensive result in Table 1, where each column shares the same question context but varies in the number of answer choices. A many-choice question contains all the options from the corresponding few-choice versions. For example, each question in the 8-choice task includes all the choices from the 6-choice task. As shown in Table 1, LLMs generally perform worse on many-choice QA than few-choice QA, This indicates the confidence distortion problem for LLM on many-choice QA.
3.2 LLM Inference with Self-Ensemble
We propose Self-Ensemble to solve the confidence distortion problem on many-choice QA. The intuition of Self-Ensemble is to divide a $K$-choice (many-choice) QA into multiple $m$-choice (few-choice) QA, where $m\!\ll\!K$. Then, it collects LLM’s answer probability in each few-choice case, and estimates the expected probability of each choice for the final decision. Specifically, Self-Ensemble has three steps as follows:
Step 1.
 Given the $K$ choices $\mathcal{C}\!=\!\{{op}_{1},\cdots\!,{op}_{K}\}$, Self-Ensemble randomly splits them into $\lceil\frac{K}{m}\rceil$ groups with seed $s$, which is given by $\mathcal{G}_{s}(\mathcal{C})=\{{G}_{1},\cdots,{G}_{\lceil\frac{K}{m}\rceil}\}$. Each group ${G}_{i}$ has $m$ choices, while the last group may have less than $m$ choices if $K\!\!\mod{m}\!\neq\!0$. Different groups take random subsets of $\mathcal{C}$ without overlaps, i.e. ${G}_{i}\cap{G}_{j}=\varnothing$ for $i\neq j$, and ${G}_{1}\cup\cdots\cup{G}_{\lceil\frac{K}{m}\rceil}=\mathcal{C}$.
Step 2.
For each group ${G}_{j}=\{\widetilde{op}_{1},\cdots,\widetilde{op}_{\tilde{m}}\}$, Self-Ensemble collects the probability of each choice within this group from LLMs as follows:
$Pr(\widetilde{op}_{1}),\cdots,Pr(\widetilde{op}_{\tilde{m}})=f([Q,{G}_{j}])$ (1)
Step 3.
To estimate the final probability of each choice $op_{i}$, Self-Ensemble averages all the probabilities of choice $op_{i}$ as follows:
$\!\!\!\!{Pr}(op_{i})\!=\!\mathbb{E}_{s\sim U(\mathbb{N}),{G}_{j}\sim\mathcal{G}_{s}(\mathcal{C}),\atop\widetilde{op}\sim{G}_{j}}\!\!\big{[}\mathbb{I}(op_{i}\!=\!\widetilde{op})Pr(\widetilde{op})\big{]}\!\!\!\!\!\vspace{-2mm}$ (2)
where $s\!\sim\!U(\mathbb{N}),{G}_{j}\!\sim\!\mathcal{G}_{s}(\mathcal{C})$ represents randomly splitting the choice set using different integer random seed $s$; and $\mathbb{I}(op_{i}\!=\!\widetilde{op})\!=\!1$ if $op_{i}\!=\!\widetilde{op}$; otherwise it takes 00. The intuition of Equation (2) is to estimate the expected probability of each choice by averaging over different group partitions.
An example of Self-Ensemble is given in Figure 4. Specifically, in each trial, Self-Ensemble simplifies the given 4-choice QA into 2-choice QAs by randomly grouping the choices, and achieves the choice probability within each group from LLMs. After $N$ trials, it estimates the expected probability of each choice by averaging different trials.
Figure 5: Plug-in Self-Ensemble: by incorporating the attention mask and positional re-encoding, LLMs can achieve the ensembled results in a single forward pass.
4 Plugging Self-Ensemble into LLMs
We propose integrating Self-Ensemble into existing LLM architectures to enable intrinsic self-ensemble processing during inference. As illustrated in Equations (1) and (2), Self-Ensemble requires multiple forward passes through the LLM, which are $f([Q,{G}_{1}])$, $f([Q,{G}_{2}])$, $f([Q,{G}_{3}]),\cdots$. For efficient self-ensemble processing, we design specialized attention masks and positional encoding, such that these multiple calls of LLM can be executed in a single forward pass over the concatenated sequence $f([Q,{G}_{1},{G}_{2},{G}_{3},\cdots])$. Here, each group ${G}_{j}\sim\mathcal{G}_{s}(\mathcal{C}),s\sim\mathbb{N}$ is sampled from the choice set following Section 3.2 Step 1; and $[Q,{G}_{1},{G}_{2},{G}_{3},\cdots]$ denotes combining the question and different groups of choice into a sequence of tokens as prompts for LLMs.
4.1 Attention Masking
The standard auto-regressive attention mask follows $\mathbb{I}(j\leq i)$, ensuring that each token attends to all preceding tokens in the sequence. However, following Self-Ensemble, the LLM should restrict its attention to a single group of choices during each trial, ignoring the influence of other groups. Therefore, for the concatenated input sequence with the question and all choice groups $[Q,{G}_{1},{G}_{2},{G}_{3},\cdots]$, when processing a certain choice group $G_{i}$, the LLM should not have attention to other groups $G_{j}$, $j\neq i$.
To implement this constraint, Self-Ensemble designs a custom attention mask to block the cross-group attention. As shown in the left-hand side of Figure 5, a token within group $G_{1}$ (highlighted in red) attends only to the question and other tokens within $G_{1}$. This ensures that the model focuses solely on the current group when generating the answer. In general, let $\texttt{Tk}_{i}$ denote a certain token at position $i$. Its attention mask to a previous token $\texttt{Tk}_{j}$, $j\leq i$ is given by
$\!\!\!\!\!M_{i,j}\!=\!\begin{cases}1,&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{if }\texttt{Tk}_{j}\in\texttt{Tk}(Q),\\ \prod\limits_{n=1}^{N}\mathbb{I}(\texttt{Tk}_{i},\texttt{Tk}_{j}\in\texttt{Tk}(G_{n}))&\!\!\!\!\text{otherwise},\!\!\!\!\!\!\end{cases}$ (3)
where $\texttt{Tk}(Q)$ denotes the tokens of question $Q$; $\mathbb{I}(\texttt{Tk}_{i},\!\texttt{Tk}_{j}\!\in\!\texttt{Tk}(G_{n}))\!=\!1$ if $\texttt{Tk}_{i}$ and $\texttt{Tk}_{j}$ belong to the same group, and 00 otherwise; and $M_{i,j}=0$ for $j>i$ as standard attention masks. This attention mask enforces pairwise independence between choice groups during inference.
4.2 Positional Re-encoding
The standard auto-regressive positional encoding follows $\texttt{Pos}_{i}\!=\!i$, where $\texttt{Pos}_{i}$ denotes the positional index of a token at position $i$. Following Self-Ensemble, the LLM should process each question-group pair $[Q,{G}_{j}]$ as a continuous input sequence for $1\!\leq\!j\!\leq\!N$. However, this continuity is disrupted in the concatenated input sequence $[Q,G_{1},G_{2},G_{3},\cdots]$, because starting from $G_{2}$, the question $Q$ and a choice group are no longer physically adjacent. For example, ${G}_{1}$ lies between $Q$ and ${G}_{2}$, breaking the positional continuity between $Q$ and $G_{2}$. To preserve the relative positional relationships between the question and each choice group, each choice group should be encoded into the relative contextual position to the question. Specifically, the positional indexes of question tokens and choice tokens are given as follows:
Question tokens.
A question token $\texttt{Tk}_{i}\!\in\!\texttt{Tk}(Q)$ should have absolute position $\texttt{Pos}_{i}=i$.
Choice tokens.
A choice token $\texttt{Tk}_{i}$, if it is with choice group $G_{n}$, its absolute position should be replaced into a relative position as follows:
$\texttt{Pos}_{i}=i-\sum_{j=1}^{n}|\texttt{Tk}(G_{j})|,$ (4)
where $|\texttt{Tk}(G_{j})|$ indicates the token number of a choice group $G_{j}$; and $n$ takes maximal value that satisfies $i\geq\sum_{j=1}^{n}|G_{j}|$.
Following Equation (4) to replace the position index of choice tokens, the position index of input sequence $[Q,{G}_{1},\cdots,{G}_{n}]$ is given by
$\Big{[}\overbrace{0,1,\cdots,|\texttt{Tk}(Q)|\!-\!1}^{\text{Position index of Q}},\!\!\!\!\!\!\!\!$ (5)
$\overbrace{|\texttt{Tk}(Q)|,\cdots,|\texttt{Tk}(Q)|\!+\!|\texttt{Tk}(G_{1})|\!-\!1}^{\text{Position index of $G_{1}$}},\cdots,$
$\overbrace{|\texttt{Tk}(Q)|,\cdots,|\texttt{Tk}(Q)|\!+\!|\texttt{Tk}(G_{n})|\!-\!1}^{\text{Position index of $G_{n}$}}\Big{]}$
By combining the attention mask with positional re-encoding, we ensure that processing the concatenated sequence $[Q,G_{1},G_{2},G_{3},\cdots]$ is functionally equivalent to independently processing each question-group pair $[Q,G_{1}],[Q,G_{2}],[Q,G_{3}],\cdots$. This enables Self-Ensemble to seamlessly integrate into existing LLM architectures.
4.3 Choice Probability
By using the attention mask and positional re-encoding in Equations (3) and (5), respectively, the LLM can simultaneously process all groups of choices in a single forward pass. As shown in Figure 5, within the output sequence, the ending token of each group captures the choice probabilities within each choice group. Formally, let $\boldsymbol{a}=f([Q,{G}_{1},{G}_{2},{G}_{3},\cdots])$ denote the sequence of output logits. For a certain choice $\widetilde{op}_{i}\in{G}_{j}$, the probability of $\widetilde{op}_{i}$ is estimated by
$Pr(\widetilde{op}_{i})=\boldsymbol{a}[\texttt{index}_{j},\texttt{index}_{i}],$ (6)
where $\texttt{index}_{j}\!=\!|Q|\!+\!\sum_{k=1}^{j-1}\!|\texttt{Tk}(k)|\!-\!1$ takes the ending token position of group $G_{j}$; $\texttt{index}_{i}$ takes the token ID of $\widetilde{op}_{i}$ defined by the LLM tokenizer. The final probability of each choice takes the expected probability $Pr({op}_{i})\!=\!\mathbb{E}[\mathbb{I}({op}_{i}\!=\!\widetilde{op}_{i})Pr(\widetilde{op}_{i})]$.
The advantage of Self-Ensemble lies in its seamless integration with existing LLM architectures. It enables the LLM to achieve the expected probability of each choice in a single forward pass, avoiding separate inference over individual choice groups. This enables Self-Ensemble to behave both accurately and efficiently on MCQA.
QASC TruthfulQA MMLU-Pro Biology Average
Model Method Accuracy Improve Accuracy Improve Accuracy Improve Accuracy Improve
LLaMA-3-8B Standard inference 73.22 - 51.99 - 62.00 - 62.40 -
Vector Scaling 75.16 +1.94 57.40 +5.41 62.00 +0.00 64.85 +2.45
Dirichlet Calibration 74.96 +1.74 53.07 +1.06 62.24 +0.24 63.42 +1.02
Self-Ensemble 80.67 +7.45 59.57 +7.58 65.78 +3.78 68.67 +6.27
Mistral-7B-v0.1 Standard inference 57.78 - 32.13 - 50.89 - 46.93 -
Vector Scaling 58.42 +0.64 41.52 +9.39 51.56 +0.67 50.50 +3.57
Dirichlet Calibration 58.53 +0.75 35.38 +3.25 51.11 +0.22 48.34 +1.41
Self-Ensemble 66.31 +8.53 54.15 +22.02 55.11 +4.22 58.52 +11.59
Qwen-2-7B Standard inference 69.44 - 53.79 - 59.78 - 61.00 -
Vector Scaling 73.00 +3.56 55.60 +1.81 60.44 +0.66 63.01 +2.01
Dirichlet Calibration 75.05 +5.61 56.68 +2.89 62.00 +2.22 64.58 +3.58
Self-Ensemble 77.86 +8.42 60.29 +6.50 66.00 +6.22 68.05 +7.05
Table 2: Accuracy of Self-Ensemble on the QASC, TruthfulQA, MMLU-Pro Biology datasets.
5 Experiment
In this experiment, we conduct experiments to evaluate Self-Ensemble, aiming to answer the following research questions: RQ1: Can Self-Ensemble improve on LLM’s accuracy on many-choice QA? RQ2: Can Self-Ensemble mitigate the confidence distortion problems for LLMs? RQ3: How does Self-Ensemble reshape the scaling of model accuracy with respect to parameter size?
5.1 Experimental Setup
We specify the datasets, LLMs, evaluation metrics, and implementation details.
Models.
We evaluate Self-Ensemble using three popular model families: LLaMA-3-8B Touvron et al. (2023), Mistral-7B-v0.1 Jiang et al. (2024), and Qwen-2-7B Yang et al. (2024). We download these models from the Huggingface Transformers library Wolf et al. (2019).
Dataset.
The evaluation of Self-Ensemble is based on the QASC Khot et al. (2020), TruthfulQA Lin et al. (2021), and MMLU-Pro Wang et al. (2024d) datasets. QASC: It is a multi-hop reasoning dataset comprised of 8-choice QA. We evaluate LLM performance on its validation set with 927 questions under a closed-book setting. TruthfulQA: It has 817 multiple-choice and short-answer questions across 38 categories, such as health, law, and finance. Following many-choice setting, we filter out questions with fewer than five incorrect choices. As a result, 277 questions have at least one correct choice and five incorrect choices. MMLU-Pro: Enhancement of the MMLU dataset by selecting questions where state-of-the-art LLMs consistently fail. We use the Biology subset for our experiments, which has 450 10-choice questions.
Baseline Methods.
Standard inference: We perform standard inference by comparing the probability of each choice token and selecting the option with the highest likelihood. Vector scaling: Vector scaling Guo et al. (2017) applies a learnable scale vector $w$ and bias $b$ to the logits before softmax, enabling finer-grained calibration. The parameters are optimized to maximize accuracy on a validation set. Dirichlet calibration: Dirichlet calibration Zong et al. (2024) is a lightweight multiclass calibration method that adjusts the model’s predicted probability to better match observed frequencies on a validation set.
Implementation Details.
For QASC’s 8-choice QA, Self-Ensemble splits each question into a 4-choice QA for 20 trials. In addition, for the TruthfulQA dataset, it splits the 6-choice QAs into 3-choice QAs for 6 trials. Moreover, for the MMLU-Pro Biology dataset, it splits the 10-choice QAs into 5-choice QAs for 40 trials. Self-Ensemble does not reply on a validation set for optimizing any parameter or special setting.
Figure 6: Probability of model confidence exceeding a threshold on the QASC, TruthfulQA, and MMLU-Pro-Biology dataset, for each model under both correct- and incorrect-answer conditions.
5.2 Accuracy of MCQA (RQ1)
Table 2 shows the accuracy (%) of Self-Ensemble. These results are compared with baseline methods and standard inference of LLMs.
Accuracy Improvement.
As shown in Table 2, Self-Ensemble consistently outperforms LLM standard inference and baseline methods, demonstrating its potential in enhancing LLM’s performance on MCQA. Moreover, compared with baseline methods, Self-Ensemble does not require a validation set for maximizing performance.
Model Agnosticism.
By integrating with different families of LLMs, Self-Ensemble shows consistent performance and improvement, as shown in Table 2. This generality indicates that Self-Ensemble can potentially serve as a versatile enhancement for a range of LLMs in practice.
Stable Improvement.
For datasets of different difficulty levels, standard inference shows less accuracy on TruthfulQA and MMLU-Pro than QASC. This implies that the confidence distortion is general problem on both easy and hard MCQA tasks. In contrast, Self-Ensemble consistently improves performance across all difficulty levels, demonstrating its general effectiveness in mitigating the confidence distortion problem.
5.3 Confidence Calibration by Self-Ensemble (RQ2)
In this section, we show that Self-Ensemble mitigates the confidence distortion problem of LLMs in Figure 6. Specifically, Figure 6 shows the proportion of $f([Q,\mathcal{C}])\!>\!\tau$ versus $\tau$ for $10\!<\!\tau\!<\!$ 1 on three different LLMs and datasets, where $f$ takes different LLMs; and $f([Q,\mathcal{C}])\!>\!\tau$ represents LLM’s predicted probability exceeds a threshold $\tau$. A higher value of $f([Q,\mathcal{C}])\!>\!\tau$ proportion across $10\!<\!\tau\!<\!$ 1 indicates strong model confidence.
In Figure 6’s sub-figures (a)-(c), (g)-(i), (m)-(o), Self-Ensemble has a higher confidence level on its correct answers; while in sub-figures (d)-(f), (j)-(l), (p)-(r), Self-Ensemble has a lower confidence level on its incorrect answers. This indicates Self-Ensemble can effective mitigate the under-confident problems on correct answers and over-confident problems on incorrect answers, improving the reliability of LLMs on MCQA in practice.
Figure 7: Qwen2.5-1.5B, 3B, 7B, and 14B with standard inference and Self-Ensemble on the QASC dataset.
5.4 Improving Accuracy-to-Parameter Scaling (RQ3)
Self-Ensemble achieves improved scaling of accuracy with parameter size. As shown in Figure 7, Self-Ensemble consistently yields higher accuracy on the QASC dataset as the model size increases from 1.5B to 14B. It enables better scalability across model sizes and can be integrated with any LLM without requiring additional data or retraining. Notably, Qwen2.5-1.5B, 3B, 7B with Self-Ensemble outperform larger models using standard inference. This demonstrates the potential of Self-Ensemble to advance the deployment of small language models in real-world scenarios with constrained memory resources.
5.5 Application to Quantized LLMs
To further explore Self-Ensemble’s application to memory-constrained settings, we show its effectiveness on quantized LLMs, including LLaMA-3-8B, Mistral-7B-v0.1, and Qwen2-7B under 4-bit bitsandbytes (BNB) quantization. As shown in Table 3, Self-Ensemble effectively offsets accuracy loss caused by the aggressive compression, enabling each quantized model to outperform its full-precision version. This indicates Self-Ensemble’s potential for quantized LLMs in resource-constrained environments, without requiring additional training or external information.
Model 4-bit Full Precision 4-bit + Self-Ensemble
LLaMA-3-8B 72.57 73.22 79.16
Mistral-7B-v0.1 57.45 57.78 65.23
Qwen2-7B 67.71 69.44 78.83
Table 3: Self-Ensemble helps quantized LLMs outperform full-precision LLMs on the QASC dataset.
Model w/o Attn Mask w/o Pos Re-enc Self-Ensemble
LLaMA-3-8B 19.22 67.39 80.67
Mistral-7B-v0.1 18.57 53.24 66.31
Qwen2-7B 15.44 71.27 77.86
Table 4: Comparison of Self-Ensemble with that w/o attention masks or positional re-encoding.
5.6 Ablation Study
We demonstrate the individual contribution of attention masks and positional re-encoding to Self-Ensemble. Specifically, we conduct experiments comparing Self-Ensemble with its ablated versions: without attention masks or positional re-encoding. Experimental results for the LLaMA-3-8B, Mistral-7B-v0.1, and Qwen2-7B on the QASC dataset are given in Table 4. It is observed that each model has a significant loss of accuracy without attention masks or positional re-encoding, indicating that both attention structure and position encoding contributes to the self-ensemble process. Self-Ensemble has a special attention mask designed in Sections 4.2 and positional re-encoding in Section 4.1 to enable intrinsic LLM inference over different choice groups and produce ensemble results efficiently. This can effectively calibrate LLM confidence and enhance the accuracy on many-choice problems.
6 Related Work
Confidence Calibration.
Calibration aligns predicted probabilities with true correctness likelihoods to ensure reliable confidence estimates. Traditional methods like vector scaling Guo et al. (2017), isotonic regression Zadrozny and Elkan (2002), and Dirichlet calibration Kull et al. (2019) have been adapted to LLMs but often assume fixed prediction heads and struggle with distribution shifts or multi-choice uncertainty. Recent generative-specific techniques, such as logit perturbation Jiang et al. (2021), prompt-based reweighting Zhou et al. (2022b), and Bayesian post-hoc methods Wang et al. (2023), partially address these limitations. However, consistent calibration across diverse LLM architectures and tasks remains challenging, highlighting the need for effective, model-agnostic solutions.
LLM Ensemble.
Ensembling has long been a powerful strategy to improve both model accuracy and robustness. In LLMs, methods such as voting schemes Zhou et al. (2022a), stochastic decoding Ho and Weller (2022), and diverse prompting Wang et al. (2022) aggregate outputs from multiple model instances or inference runs. These methods are especially useful for reducing variance in generative outputs and mitigating individual model biases. Additionally, ensemble approaches support calibration via confidence averaging and self-consistency mechanisms Wang et al. (2022). However, standard ensembles are computationally costly and less generalizable across model families and tasks, motivating for a more efficient and scalable ensemble method.
7 Conclusion
In this work, we demonstrate the confidence distortion problem of LLMs generally on MCQA, particularly in the many-choice setting. To solve this problem, we propose Self-Ensemble by decomposing a many-choice problem into several few-choice problems, and aggregating the intermediate results into the final solution. We further integrate Self-Ensemble with existing LLM architecture, enabling an intrinsic self-ensemble process for LLM inference. Experimental results across various LLMs and datasets show that Self-Ensemble can effectively overcome this problem, improving confidence in correct answers and reducing confidence in incorrect answers. This enables Self-Ensemble to significantly improve the accuracy of LLMs in MCQA by $8\%$ on average.
8 Limitations and Potential Risks
In this work, we propose a Self-Ensemble to calibrate the LLM decisions on many-choice QA. The application of Self-Ensemble is limited to the muti-choice question-answer (MCQA) problems. While MCQA stands out as a standard and challenging benchmark to evaluate the ability of LLMs, there are open-ended QA tasks in real-world scenarios. Calibrating LLMs’ decisions on open-ended QA remains our future research. Furthermore, Self-Ensemble is methodologically orthogonal to existing calibration techniques such as Vector Scaling and Dirichlet Calibration, enabling synergistic integration to further enhance performance.
Appendix
Appendix A General-purpose LLMs with Self-Ensemble Outperforms Reasoning LLMs
We show that general-purpose LLMs, when combined with Self-Ensemble, can surpass specialized reasoning LLMs on MCQA. Deepseek-R1-Distill models Guo et al. (2025) are distilled LLMs that internalize reasoning by fine-tuning on retrieved traces from Deepseek-R1. As shown in Table 5, both LLaMA-3-8B and Qwen2-7B, after applying Self-Ensemble, achieve higher accuracy compared to their reasoning-tuned counterparts Deepseek-R1-Distill-Llama-8B and Deepseek-R1-Distill-Qwen-7B on the QASC dataset. This highlights that Self-Ensemble can unlock the potential of general LLMs on MCQA without requiring dedicated reasoning fine-tuning.
Appendix B Self-Ensemble Outperforms Chain of Thought Prompting
We demonstrate that Self-Ensemble yields higher improvement on the QASC dataset than chain of thought (CoT) prompting Wei et al. (2022). CoT elicits intermediate reasoning steps from an LLM by prepending a prompt such as “Let’s think step by step” before generating the final answer to improve accuracy on MCQA. As shown in Table 5, with Self-Ensemble, both LLaMA-3-8B and Qwen2-7B models achieve higher accuracy on the QASC dataset than CoT prompting. While CoT prompting incurs additional inference latency and token-generation overhead by producing reasoning chains, Self-Ensemble shows its potential as a more efficient and effective calibration method on MCQA, requiring only a single forward pass with marginal overhead.
Appendix C Packages
In this work, we use the transformers along with datasets packages Wolf et al. (2019) for model and dataset loading. All open-sourced packages have the Apache-2.0 license, which allows for academic research.
Model Accuracy
LLaMA-3-8B 73.22
DeepSeek-R1-Distill-Llama-8B 73.33
LLaMA-3-8B with CoT 74.62
LLaMA-3-8B with Self-Ensemble 80.67
Qwen2-7B 69.44
DeepSeek-R1-Distill-Qwen-7B 70.63
Qwen2-7B with CoT 74.19
Qwen2-7B with Self-Ensemble 77.86
Table 5: General LLMs with Self-Ensemble outperform reasoning LLMs and chain of thought prompting on the QASC dataset.
Name Value
Data type torch.bfloat16
Flash-Attention False
Eval batch-size 1
Computing Infrastructure GPU
GPU Model NVIDIA-A40
GPU Memory 48GB
GPU Number 1
CUDA Version 12.3
CPU Memory 512GB
Table 6: Experiment configuration and computing infrastructure.
Appendix D Computational Infrastructure
The computational infrastructure information is given in Table 6.